# Learning in Chaos: Efficient Autoscaling and Self-healing for Distributed Training at the Edge

### v1

@@ III Multi-neighbor State Replication
When a new node joins, it needs to pull training states from existing nodes, including model weights, optimizer states, and runtime info like epoch, iteration, hyperparameters, and so on. These states are consistent across nodes because the training is synchronous. As shown in Fig. 3, model weights and optimizer states take up most of the size. Even pulling from a neighbor can take up to one minute, which is far too slow for an edge cluster with frequent node joins. To speed things up, this section models a shard assignment problem based on multi-neighbor replication, and propose a fast algorithm to cut the scale-out delay.
Figure 3: Overview of replication delays for each component. The link bandwidth is set to 200 Mbps.
III-A Problem Model
Let $G=(V,E)$ be a graph shown in Fig. 1a, with each edge has a weight representing the delay of transmitting one unit of data over that link. Given a DML task with training state $w$ (e.g., model weights and optimizer states). For a fine-grained control over traffic load, we split $w$ into $K$ shards, each with data size $s$. Assume $s$ can divide the total data size $|w|$, that is, $|w|=Ks$. Let $\mathcal{K}=\{w_{1},w_{2},\cdots,w_{K}\}$ be the set of model shards. Now, a new node $v_{\text{new}}$ joins and needs to pull shards $\mathcal{K}=\{\mathcal{K}_{1},\cdots,\mathcal{K}_{|\mathcal{U}|}\}$ from its neighbors $\mathcal{U}$, where each neighbor $u\in\mathcal{U}$ provides a shard subset $\mathcal{K}_{u}\subset\mathcal{K}$. These shard subsets are disjoint and their collection covers all shards, i.e., $\bigcup_{u\in\mathcal{U}}\mathcal{K}_{u}=\mathcal{K}$ and $\mathcal{K}_{u}\cap\mathcal{K}_{v}=\emptyset$ for any $u\neq v$.
Let $t^{\text{prop}}_{u\rightarrow v_{\text{new}}}$ be the propagation delay and $t^{\text{trans}}_{u\rightarrow v_{\text{new}}}$ the per-parameter transmission delay from neighbor $u$ to $v_{\text{new}}$. Then, the delay to send all shards in the set $\mathcal{K}_{u}$ is:
@@ III-A Problem Model
As shown in Fig. 2b, state replication starts after the all-reduce, but neighbors finish all-reduce at different times. Let $\tau_{u}^{\text{sync}}$ be the time when neighbor $u$ finishes all-reduce. Our goal is to minimize the latest completion time any neighbor finishes its replication:
$\min_{s,\mathcal{K}_{u}}\max_{u\in\mathcal{U}}\quad$ $t_{u}(s,\mathcal{K}_{u})+\tau_{u}^{\text{sync}},$ (2)
s.t. $s\in\mathbb{Z}^{+},|\mathcal{K}_{u}|>0,\forall u\in\mathcal{U}.$ (3)
To formally represent $\mathcal{K}_{u}$, we introduce a binary selection matrix $x_{uj}\in\{0,1\}$, where $u\in\mathcal{U}$ is a neighbor of the new node and $j\in\{1,2,\cdots,K\}$ indexes the model shards. If $x_{uj}=1$, neighbor $u$ sends shard $w_{j}$; otherwise, it doesn’t. Thus, the problem can be reformulated as P1:
$\min_{s,x_{uj}}\max_{u\in\mathcal{U}}\quad$ $t_{u}(s,x_{uj})+\tau_{u}^{\text{sync}},$ (4)
s.t. $t_{u}(s,x_{uj})=t^{\text{prop}}_{u\rightarrow v_{\text{new}}}+\sum_{j=1}^{K}s\cdot t^{\text{trans}}_{u\rightarrow v_{\text{new}}}\cdot x_{uj},$ (5)
@@
$s\in\mathbb{Z}^{+},|\mathcal{K}_{u}|>0~{}\forall u\in\mathcal{U}.$ (7)
This objective involves a max operation, so we introduce an auxiliary variable $\theta$ to bound the completion time across all neighbors. Then the problem becomes P2:
$\min_{s,x_{uj},\theta}\quad$ $\theta$ (8)
@@
$s\in\mathbb{Z}^{+},x_{uj}\in\{0,1\}~{}\forall u,j.$ (12)
The constraint includes a product of variables, $s\cdot x_{uj}$, where $s$ is an integer and $x_{uj}$ is binary. This makes the problem a mixed integer nonlinear programming (MINLP) and hard to solve. To make it easier, we hope to fix $s$ in P2 so that it becomes a linear program.
If we fix $x_{uj}$ and optimize only $s$, the problem P2 becomes:
$\min_{s}\max_{u\in\mathcal{U}}~{}\alpha_{u}\cdot s+\beta_{u},~{}\text{s.t.}~{}s\in\mathbb{Z}^{+},$ (13)
where $\alpha_{u}>0,\beta_{u}>0$ are constants. Since the objective is a max over linear functions, the overall objective is monotonically increasing with respect to $s$. Although $\theta$ in P2 is not strictly monotonic, it often exhibits a quasi-monotonic trend in practice, making binary search an effective way to quickly find a near-optimal $s$, as described in Algorithm 1.
Then, we can apply an external binary search to fix $s$ and focus on optimizing $x_{uj}$. The problem P2 now becomes P3:
@@
$x_{uj}\in\{0,1\},~{}\forall u,j.$ (18)
P3 is a mixed integer linear programming (MILP) with binary variables. It is more tractable to solve than a MINLP in P2, with standard MILP solvers. By iterating over possible values of $s$ using binary search, we can efficiently find the optimal shard assignment $\mathcal{K}_{u}^{*},~{}\forall u\in\mathcal{U}$.
However, even as a standard MILP, P3 remains NP-hard. With large models and complex networks, solving P3 can take several seconds. Since P2 uses binary search over $s$, P3 must be solved multiple times, pushing the total delay into tens of seconds. This solution delay even exceeds the state replication time, making the MILP solver impractical. Worse, every new node triggers a fresh solve, further tightening the latency constraints. To keep things fast, we trade the mathematically elegant but slow MILP solver for a fast greedy algorithm (Algorithm 2). This greedy solution provides a bounded optimality gap and can find a near-optimal solution in a flash, making it more well-suited for practical use.
Input: Training states $w$, neighbor set $\mathcal{U}$, new node $v_{\text{new}}$.
Output: Shard assignment $\{\mathcal{K}_{u}^{*},~{}\forall u\in\mathcal{U}\}$, Shard size $s^{*}$.
1
2Initialize $s_{l}\leftarrow\text{min\_layer\_size}(w)$, $s_{h}\leftarrow\text{max\_layer\_size}(w)$;
3Initialize the objective value $\theta^{*}\leftarrow+\infty$;
4while $s_{l}\leq s_{h}$ do
5       Set $s\leftarrow\lfloor\frac{s_{l}+s_{h}}{2}\rfloor$;
6      Split tensors in $w$ into a set of shards $\mathcal{K}$, each of size $s$;
7      Solve P3 with given $\mathcal{K}$ and $s$:
8       $\{\mathcal{K}_{u},\forall u\}\leftarrow\mathsf{Greedy\ Shard\ Assignment}(\mathcal{K},\mathcal{U},s,v_{\text{new}})$;
9      Calculate objective value $\theta$ in P2 with given $\mathcal{K}_{u}$ and $s$;
10      if $\theta<\theta^{*}$ then
11             $\theta^{*}\leftarrow\theta$, $\mathcal{K}_{u}^{*}\leftarrow\mathcal{K}_{u}$, $s^{*}\leftarrow s$,
12             $s_{h}\leftarrow s-1$;
13
14      else
15             $s_{l}\leftarrow s+1$;
16
17
return $\{\mathcal{K}_{u}^{*},~{}\forall u\in\mathcal{U}\},s^{*}$;
Algorithm 1 $\mathsf{Shard\ Assignment\ with\ Binary~{}Search~{}(P2)}$
Input: Model shard set $\mathcal{K}$, neighbor set $\mathcal{U}$, shard size $s$, new node $v_{\text{new}}$, propagation delay $t_{u\rightarrow v_{\text{new}}}^{\text{prop}}$, transmission delay $t_{u\rightarrow v_{\text{new}}}^{\text{trans}}$, synchronization delay $\tau_{u}^{\text{sync}}$ for each $u\in\mathcal{U}$,
@@ 15             $s_{l}\leftarrow s+1$;
2Initialize the assignment $\mathcal{K}_{u}\leftarrow\emptyset,~{}\forall u\in\mathcal{U}$;
3
4Initialize the load $l_{u}=t_{u\rightarrow v_{\text{new}}}^{\text{prop}}+\tau_{u}^{\text{sync}},~{}\forall u\in\mathcal{U}$;
5foreach shard $w_{k}\in\mathcal{K}$ of size $s$ do
6       Select the neighbor $u^{*}$ with least estimated load: $u^{*}\leftarrow\arg\min_{u}(l_{u}+s\cdot t_{u\rightarrow v_{\text{new}}}^{\text{trans}})$;
7
8      Assign the shard $w_{k}$ to neighbor $u^{*}$: $\mathcal{K}_{u^{*}}\leftarrow\mathcal{K}_{u^{*}}\cup\{w_{k}\}$;
9       Update the load: $l_{u^{*}}\leftarrow l_{u^{*}}+s\cdot t_{u^{*}\rightarrow v_{\text{new}}}^{\text{trans}}$;
10
11
@@
Algorithm 2 $\mathsf{Greedy\ Shard\ Assignment~{}(P3)}$
@@ III-B Algorithm Design
The training state $w$ contains many tensors of varying sizes, small ones like convolution kernels and biases, and large ones like fully connected layer weights. Assigning these raw tensors directly to neighbors creates uneven traffic loads, causing the coflow completion time to spike. To better balance traffic, we introduce $s$ in P2 to control the model shard size after splitting. If $s$ is set too large, shards are few and traffic stays uneven; but if it is set small, transmission becomes inefficient. Leveraging the quasi-monotonic nature of P2 with respect to $s$, Algorithm 1 uses binary search to quickly find a good value. In each iteration, it splits the training state $w$ into shards $\mathcal{K}=\{w_{1},w_{2},\cdots,w_{K}\}$ of size $s$, then calls the P3 solver (Algorithm 2) to assign shards to neighbors. If the result improves, binary search continues on a smaller size; otherwise, it tries a larger size. This search repeats until the optimal shard size $s^{*}$ and assignment $\{\mathcal{K}_{u}^{*},~{}\forall u\in\mathcal{U}\}$ are found.
To solve P3, Algorithm 2 takes a greedy approach: it iteratively assigns shards to the neighbor with the lowest estimated load. To see how the estimate load $l_{u}$ is updated, we rewrite P3 as the following min-max problem:
$\min_{x_{uj}}\max_{u\in\mathcal{U}}\quad\underbrace{t^{\text{prop}}_{u\rightarrow v_{\text{new}}}+\tau_{u}^{\text{sync}}}_{\text{initial term}}+\sum_{j=1}^{K}\underbrace{s\cdot t^{\text{trans}}_{u\rightarrow v_{\text{new}}}}_{\text{update term}}\cdot x_{uj}$ (19)
Following this pattern, we initialize $l_{u}$ using the fixed term (the initial term), $t^{\text{prop}}_{u\rightarrow v_{\text{new}}}+\tau_{u}^{\text{sync}}$, and increment it with the per-shard cost (the update term), $s\cdot t^{\text{trans}}_{u\rightarrow v_{\text{new}}}$. The initial term captures the base delay before neighbor $u$ can start transmitting a shard, and the update term captures the delay for transmitting one shard to the new node. Then, it looks at all neighbors, picks the neighbor $u^{*}$ with the lowest estimate load, assigns the shard to $u^{*}$, and updates its load $l_{u^{*}}$ accordingly.
Optimality Gap. This greedy solution tries to equalize the delay by always choosing the least-loaded neighbor. It reduces to the longest processing time first (LPT) heuristic for the $P\|C_{\text{max}}$ scheduling problem. By Graham’s bound [24], LPT achieves a makespan no worse than $(\frac{4}{3}-\frac{1}{3|\mathcal{U}|})\cdot\text{OPT}$, where OPT is the optimal value. That means Algorithm 2 is guaranteed to stay within 33% of the optimum, and in typical cases where neighbors are few (e.g., $|\mathcal{U}|=2$), the optimality gap improves to just 17%.
Complexity Analysis. Algorithm 1 performs a binary search over the shard size $s$, with at most $O(\log S)$ steps, where $S$ is the range of shard sizes. In each step, it runs Algorithm 2 to assign $K$ shards to $|\mathcal{U}|$ neighbors, taking $O(K\cdot|\mathcal{U}|)$ per step. Thus, the total time complexity of Algorithm 1 is $O(K\cdot|\mathcal{U}|\cdot\log S)$, which is in polynomial time.
To sum up, our algorithms come very close to optimal and run orders of magnitude faster than solving the MINLP/MILP, which is further validated in our experiments.
IV Cluster Monitor and Peer Negotiation
In the previous section, we showed how shard scheduling can speed up state replication when new nodes join, but two issues remain:
  1. 1.
    The algorithms rely on prior knowledge of the cluster topology and resources, like which nodes are neighbors, and the link latency and bandwidth between them.
  2. 2.
    The algorithms are triggered by cluster changes, such as node joins, exits, and failures, so they rely on timely detection of these events to respond to changes.
In this section, we first introduce a cluster monitor to track resources for scheduling decisions and to detect node and link changes, followed by peer-negotiation protocols to handle scale-in/out and connect/disconnect-link events.
IV-A Cluster Monitor
Cluster Topology Monitoring. The primary task of the cluster monitor is to detect the overlay topology. In Chaos, this is handled by the scheduler, which tracks each node’s ID, state (e.g., active, failed, or standby), neighbors, and join time. This data helps scheduling algorithms (e.g., Algorithms 1 and 2) identify neighbors $\mathcal{U}$ and quickly spot changes to trigger scaling events. It also supports DML systems like NetStorm [25] to build their all-reduce strategies.
@@ IV-A Cluster Monitor
Scale-out Add a node, connect it to neighbors, and synchronize training state. Node joins 0.5-5 s
Scale-in Remove a node and terminate its sockets to neighbors. Node exits or failure <1 ms
Connect-link Establish a new socket between two nodes. Link joins <1 ms
Disconnect-link Terminate the socket between two nodes. Link exits or failure <1 ms

### v2

@@ III Multi-neighbor State Replication
To minimize the state replication delay when a new node joins, this section formulates a model sharding and assignment problem. The original mixed-integer nonlinear program (MINLP) P2 is intractable, so we analyze the monotonicity on a divisibility chain and use its non-decreasing nature to reduce P2 to a 0-1 mixed-integer linear program (0-1 MILP) P3. However, solving P3 with modern MILP solvers remains slow and prone to OOM. We therefore prove that P3 is a special case of MILP, equivalent to a multi-way merge problem, where a simple greedy rule yields the optimal solution. Based on this insight, we design a polynomial-time greedy algorithm and confirm its optimality through experiments.
III-A Model Sharding and Assignment Problem
The training state contains tensors of varying sizes, from small ones like convolution kernels and biases to large fully connected weights. Assigning these raw tensors directly to neighbors would cause uneven traffic. To balance the load, we split tensors into equal-sized shards and assign them across neighbors. Here, “assignment” indicates that each shard is sent by a different neighbor to the new node. If shards are too large, there would be fewer shards and traffic remains uneven; if too small, shard management becomes expensive.
Let $G=(V,E)$ be a weighted graph, with each edge having a weight representing the delay of transmitting one unit of data over that link. Given a DML task with model weights $w$, to achieve a fine-grained control over traffic load, we split $w$ into $K$ shards, each with data size $s$. Assume $s$ can divide the total data size $|w|$, that is, $|w|=Ks$. Let $\mathcal{K}=\{w_{1},w_{2},\cdots,w_{K}\}$ be the set of model shards. Now, a new node $v_{\text{new}}$ joins and needs to fetch shards $\mathcal{K}=\{\mathcal{K}_{1},\cdots,\mathcal{K}_{|\mathcal{U}|}\}$ from its neighbors $\mathcal{U}$, where each neighbor $u\in\mathcal{U}$ is assigned an non-empty shard subset $\mathcal{K}_{u}\neq\emptyset,\mathcal{K}_{u}\subset\mathcal{K}$. These shard subsets are disjoint and their collection covers all shards, i.e., $\bigcup_{u\in\mathcal{U}}\mathcal{K}_{u}=\mathcal{K}$ and $\mathcal{K}_{u}\cap\mathcal{K}_{v}=\emptyset$ for any $u\neq v$.
Let $t^{\text{prop}}_{u\rightarrow v_{\text{new}}}$ be the propagation delay and $t^{\text{trans}}_{u\rightarrow v_{\text{new}}}$ the per-parameter transmission delay from neighbor $u$ to $v_{\text{new}}$. Then, the delay to fetch all shards in the set $\mathcal{K}_{u}$ is:
@@ III-A Model Sharding and Assignment Problem
As shown in Fig. 3b, state replication starts after the all-reduce, but neighbors finish all-reduce at different times. Let $\tau_{u}^{\text{sync}}$ be the time when neighbor $u$ finishes all-reduce. Our goal is to minimize the latest completion time among all neighbors finishing their state replication to the new node:
$\min_{s,K,\mathcal{K}_{u}}\max_{u\in\mathcal{U}}\quad$ $t_{u}(s,\mathcal{K}_{u})+\tau_{u}^{\text{sync}},$ (2)
s.t. $s\in\mathbb{Z}^{+},|\mathcal{K}_{u}|>0,\forall u\in\mathcal{U},$ (3)
$K\geq|\mathcal{U}|,~|w|=Ks.$ (4)
To formally represent $\mathcal{K}_{u}$, we introduce a binary selection matrix $x_{uj}\in\{0,1\}$, where $u\in\mathcal{U}$ is a neighbor of the new node and $j\in\{1,2,\cdots,K\}$ indexes the model shards. If $x_{uj}=1$, neighbor $u$ sends shard $w_{j}$ to the new node. Thus, the problem can be reformulated as P1:
$\min_{s,K,x_{uj}}\max_{u\in\mathcal{U}}\quad$ $t_{u}(s,x_{uj})+\tau_{u}^{\text{sync}},$ (5)
s.t. $t_{u}(s,x_{uj})=t^{\text{prop}}_{u\rightarrow v_{\text{new}}}+\sum_{j=1}^{K}st^{\text{trans}}_{u\rightarrow v_{\text{new}}}x_{uj},$ (6)
@@
$\sum_{j=1}^{K}x_{uj}\geq 1,~\forall u\in\mathcal{U},$ (8)
$x_{uj}\in\{0,1\},~\forall u\in\mathcal{U},j\in\{1,2,\cdots,K\},$ (9)
$s\in\mathbb{Z}^{+},K\geq|\mathcal{U}|,~|w|=Ks.$ (10)
This objective involves a max operation, so we introduce an auxiliary variable $\theta$ to bound the completion time across all neighbors. Then the problem is rewritten to P2:
$\min_{s,K,x_{uj},\theta}\quad$ $\theta$ (11)
@@
$\sum_{j=1}^{K}x_{uj}\geq 1,~\forall u\in\mathcal{U},$ (15)
$x_{uj}\in\{0,1\},~\forall u\in\mathcal{U},j\in\{1,2,\cdots,K\},$ (16)
$s\in\mathbb{Z}^{+},K\geq|\mathcal{U}|,~|w|=Ks.$ (17)
If both $s,K$ are treated as decision variables, the constraints contain nonlinear terms such as $sx_{uj}$ and $Ks$, with $s,K$ being positive integers and $x_{uj}$ binary. This makes P2 a mixed-integer nonlinear program (MINLP). In addition, because the summation indices and constraint sets depend on $K$, the model is not of fixed dimension, making it intractable.
To make the problem solvable, we apply an outer enumeration over $d$ to fix $s=2^{d}$ in P2, so that each subproblem reduces to a linear 0-1 program with fixed dimension.
Assumption 1.
For any neighbor $u$, its traffic completion time depends only on the total size of its assigned shards, not on $s$ or $K$ separately. In other words, we ignore per-shard extra cost (e.g., packetization, handshake), since small shards can be batched into a single data stream for transmission.
Theorem 1 (Monotonicity on a divisibility chain).
Let Assumption 1 hold, if $s_{1}\mid s_{2}$ with $s_{1}<s_{2}$, then $\theta^{(s_{1})}\ \leq\ \theta^{(s_{2})}$.
Proof.
Let $q:=s_{2}/s_{1}\in\mathbb{Z}_{>0}$. Take any feasible assignment $x^{(2)}=\{x_{uj}^{(2)}\}$ for $P2(s_{2})$ and let $n_{u}^{(2)}:=\sum_{j=1}^{K(s_{2})}x_{uj}^{(2)}$ be the number of shards assigned to neighbor $u$ under $x^{(2)}$. Then, construct $x^{(1)}=\{x_{u\ell}^{(1)}\}$ for $P2(s_{1})$ by splitting each $s_{2}$-shard into $q$ shards of size $s_{1}$, and assigning all of them to the same neighbor as before. Then for every neighbor $u$:
$n_{u}^{(1)}=q\,n_{u}^{(2)},~s_{1}\,n_{u}^{(1)}=s_{1}q\,n_{u}^{(2)}=s_{2}\,n_{u}^{(2)}.$
Hence, each neighbor’s completion time is preserved:
$t_{u}^{\text{prop}}+\tau_{u}^{\text{sync}}+s_{1}t_{u}^{\text{trans}}n_{u}^{(1)}=t_{u}^{\text{prop}}+\tau_{u}^{\text{sync}}+s_{2}t_{u}^{\text{trans}}n_{u}^{(2)}.$
Coverage is maintained since $K(s_{1})=qK(s_{2})$ and every $s_{2}$-shard is replaced by $q$ shards of size $s_{1}$. Since each neighbor holds at least one shard, $n_{u}^{(1)}=qn_{u}^{(2)}\geq 1$ whenever $n_{u}^{(2)}\geq 1$. Therefore, $x^{(1)}$ is feasible for $P2(s_{1})$ and attains the same objective value as $x^{(2)}$, which implies $\theta^{(s_{1})}\ \leq\ \theta^{(s_{2})}$. ∎
Corollary 1.
If the candidate set $\mathcal{S}=\{s_{1}<s_{2}<\cdots<s_{m}\}$ satisfies $s_{i}\mid s_{i+1}$ for all $i$, then
$\theta^{(s_{1})}\ \leq\ \theta^{(s_{2})}\ \leq\ \cdots\ \leq\ \theta^{(s_{m})},$
i.e., $\theta^{(s)}$ is non-decreasing along the divisibility chain.
According to Corollary 1, the optimal setup is $s_{1}$, which could be as small as one element or an atomic size. In practice, over-sharding (e.g., shards of size one) should be avoided, so an appropriate $s_{\min}$ should be set as the atomic size. A practical choice is the size of the smallest vector, such as biases or LayerNorm parameters. Take the Transformer model as an example: these vectors have length $H$, while most weight matrices are shaped $H\times mH$, where $m$ is an integer. Setting $s_{\min}=H$ ensures most tensors are divisible, prevents uneven shards, simplifies shard management, and still provides fine-grained load balancing across neighbors.
Thus, with $s$ and $K$ fixed, problem P2 reduces to P3:
@@
$\sum_{j=1}^{K}x_{uj}\geq 1,~\forall u\in\mathcal{U},$ (22)
$x_{uj}\in\{0,1\},~\forall u\in\mathcal{U},j\in\{1,2,\cdots,K\}.$ (23)
P3 is a standard 0-1 MILP solvable by modern solvers, but for large models with $K\in(10^{6},10^{8})$ shards and complex networks (many neighbors $|\mathcal{U}|$), solving it can still take a long time or even run out of memory, because it has $K|\mathcal{U}|+1$ decision variables. Each time a new node joins, the solver must be rerun, so millisecond-level latency is required each run. Most MILP solvers cannot meet such strict time and memory demands. Thus, a fast solver is needed.
Input: Model shard set $\mathcal{K}$, neighbor set $\mathcal{U}$, shard size $s$, new node $v_{\text{new}}$, propagation delay $t_{u\rightarrow v_{\text{new}}}^{\text{prop}}$, transmission delay $t_{u\rightarrow v_{\text{new}}}^{\text{trans}}$, synchronization delay $\tau_{u}^{\text{sync}}$ for each $u\in\mathcal{U}$.
@@
2Initialize $\mathcal{K}_{u}\leftarrow\emptyset$, $r_{u}\leftarrow t_{u\rightarrow v_{\text{new}}}^{\text{prop}}+\tau_{u}^{\text{sync}}$, $p_{u}\leftarrow s\cdot t_{u\rightarrow v_{\text{new}}}^{\text{trans}}$, and $l_{u}\leftarrow r_{u}$;
3foreach neighbor $u\in\mathcal{U}$ do
4  Pop a shard $w_{k}$ from $\mathcal{K}$ and append it to $\mathcal{K}_{u}$;
5 Update makespan $l_{u}\leftarrow l_{u}+p_{u}$;
6
7while $\mathcal{K}$ not empty do
8  Select the neighbor $u$ with least next-finish time: $u\leftarrow\arg\min_{u\in\mathcal{U}}(l_{u}+p_{u})$;
9
10 Pop a shard $w_{k}$ from $\mathcal{K}$ and append it to $\mathcal{K}_{u}$;
11 Update makespan $l_{u}\leftarrow l_{u}+p_{u}$;
12
@@ 11 Update makespan $l_{u}\leftarrow l_{u}+p_{u}$;
Algorithm 1 $\mathsf{Greedy\ Sharding\ and\ Assignment}$
@@ III-B Algorithm Design
Through an equivalent transformation, we find that P3 is equivalent to a multi-way merge problem with sorted sequences, where a greedy solution is optimal. Based on the mathematical structure of the problem, we design a greedy sharding and assignment algorithm. Both theory and experiments confirm that it yields optimal solutions and can be solved in polynomial time.
Proposition 1 (Problem Equivalence).
Let $s,K$ be fixed and constants $r_{u}=t^{\text{prop}}_{u}+\tau_{u}^{\text{sync}},p_{u}=st^{\text{trans}}_{u}$, and define aggregated counts
$n_{u}=\sum_{j=1}^{K}x_{uj},~\sum_{u\in\mathcal{U}}n_{u}=K,$ (24)
Then P3 is equivalent to P4:
$\min_{\{n_{u}|n_{u}\geq 1\}}\ \max_{u\in\mathcal{U}}\ \bigl{(}r_{u}+p_{u}n_{u}\bigr{)}\quad\text{s.t.}\quad\sum_{u\in\mathcal{U}}n_{u}=K,$ (25)
which is a multi-way merge problem with sorted sequences.
Proof.
Assigning one more shard to neighbor $u$ raises its completion time to the next value in the arithmetic progression
$\mathcal{T}_{u}=\{r_{u}+p_{u},\ r_{u}+2p_{u},\ r_{u}+3p_{u},\ldots\}.$
To realize the $j$-th completion at time $r_{u}+jp_{u}$, the 1st through (j1)(j-1)-th completions must have occurred earlier at times $r_{u}+p_{u},\ldots,r_{u}+(j-1)p_{u}$. If neighbor $u$ receives $n_{u}$ shards, the set of its completion times must be the prefix of length $n_{u}$ of $\mathcal{T}_{u}$. Hence, any feasible allocation is the same as selecting exactly $K$ elements from the union $\bigcup_{u\in\mathcal{U}}\mathcal{T}_{u}$, with at least one element from each sequence. In this way, the overall makespan equals our goal:
$\max\{\text{the chosen }K\text{ elements}\}=\max_{u\in\mathcal{U}}(r_{u}+p_{u}n_{u}).$
Then the problem becomes: choose $K$ elements from the ordered sequences $\{\mathcal{T}_{u},~\forall u\in\mathcal{U}\}$ (respecting prefix and $n_{u}\!\geq\!1$) to minimize the maximum selected value. This is exactly a multi-way merge problem with sorted sequences. ∎
Lemma 1.
The greedy method provides the optimal solution to the multi-way merge problem with sorted sequences.
Proof.
At any step, the best choice is exactly the current head of the sequences (its prefixes already taken). The greedy rule always picks the smallest head. Hence, the first $K$ outputs are the $K$ smallest feasible elements and their maximum is minimal. For example, the min-heap (a heap-based greedy) is the optimal solution to such problems [38]. ∎
Figure 4: Comparison of our greedy solution with the optimal solution on 3 models under 3–5 neighbor nodes.
@@
Scale-out Add a node, connect it to neighbors, and synchronize training state. Node joins 500ms-5s
Scale-in Remove a node and terminate its sockets to neighbors. Node exits or failure 20 ms
Connect-link Establish a new socket between two nodes. Link joins 20 ms
Disconnect-link Terminate the socket between two nodes. Link exits or failure 20 ms
Proposition 1 and Lemma 1 motivate us to design a simple but optimal greedy algorithm to solve P3:
  1. 1.
    Take the first element from each $\mathcal{T}_{u}$ to ensure $n_{u}\geq 1$;
  2. 2.
    Let $\mathcal{H}$ be the set of currently available heads;
  3. 3.
    Pick $u=\arg\min_{u\in{\mathcal{U}}}\mathcal{H}$ and update $\mathcal{H}$, repeat until $K$ elements are chosen.
In Algorithm 1, $r_{u}$ is the one-time propagation and synchronization delay from neighbor $u$, and $p_{u}$ is the per-shard transmission cost. The variable $l_{u}$ tracks the current finish time of $u$. First, each neighbor receives one shard to enforce $n_{u}\geq 1$. Then, for each remaining shard, we select the neighbor with the smallest next-finish time and assign it a shard. This greedy rule is equivalent to choosing the smallest available head in the multi-way merge view. We repeat this greedy rule until all $K$ shards are assigned.
Greedy Optimality. According to Proposition 1 and Lemma 1, the greedy rule in Algorithm 1 should produce the optimal solution. Our preliminary experiment confirms this. As shown in Fig. 4, brute-force search for P4 finds the same minimum.
Complexity Analysis. Algorithm 1 scans all $K$ shards and assigns each shard to the least-loaded neighbor among the $m=|\mathcal{U}|$ candidates, giving complexity $O(mK)$. Given a common setting with $s=2048$, $m=4$, and AlexNet, the solving time is 160ms. However, as noted in Section IV-C, this latency can be fully overlapped by all-reduce. With a min-heap, the complexity can be reduced to $O((K-m)\log m)$.
Although no new algorithm is proposed, this section shows how an intractable MINLP can be elegantly reduced to a special MILP solvable by a polynomial-time greedy algorithm that yields exact optimal solutions. This has been a significant and practical contribution.
IV Cluster Monitor and Peer Negotiation
In the previous section, we showed how model sharding and assignment can speed up state replication when new nodes join, but two issues remain:
  1. 1.
    The algorithm relies on prior knowledge of the cluster topology and resources, like which nodes are neighbors, and the link latency, bandwidth between them.
  2. 2.
    The algorithm is triggered by cluster changes, such as node joins, exits, and failures, so it relies on timely detection of these events to respond to changes.
In this section, we first introduce a cluster monitor to track resources for scheduling decisions and to detect node and link changes, followed by peer-negotiation protocols to handle self-governed scale-in/out and connect/disconnect-link events.
IV-A Cluster Monitor
Cluster Topology Monitoring. The primary task of the cluster monitor is to detect the overlay topology (an example is shown in Fig. 2a). In Chaos, this is handled by a scheduler, which tracks each node’s ID, state (e.g., active, failed, or standby), neighbors, and join time. This helps Algorithm 1 identify neighbors $\mathcal{U}$ and spot changes to trigger scaling events.